OpenStudy (anonymous):

A= 4 -2 -1 5 λ=3, 6 Find the eigenvector that corresponds to each eigenvalue

OpenStudy (anonymous):

Well I know we have to use this equation: (A−λI)v=0 Where A is the matrix and λ is an eigenvalue and I is the identity matrix but I get lost in the last step.

OpenStudy (amistre64):

|a-L b | rref = |1 b/(a-L)| | c d-L| |0 0 | ev = <-b/(a-L),1>

OpenStudy (amistre64):

since evs aint unique; we can also write it as: <-b, a-L> <b, L-a> <d-L, -c> <L-d, c> etc .....

OpenStudy (amistre64):

you seem to have frozen up :)

OpenStudy (anonymous):

I dont understand your notation

OpenStudy (anonymous):

\[(\left[\begin{matrix}4 & -2 \\ -1 & 5\end{matrix}\right]-\left[\begin{matrix}3 & 0 \\ 0 & 3\end{matrix}\right])v=0\]

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & -2 \\ -1 & 2\end{matrix}\right]v=0\]

OpenStudy (amistre64):

i just row reduced your generic 2x2 matrix with arbitrary eugene values

OpenStudy (anonymous):

\[v=\left(\begin{matrix}v1 \\ v2\end{matrix}\right)\] v1 - 2*v2=0 -v1+ 2*v2=0 What do I do from here

OpenStudy (amistre64):

row reduce the matrix

OpenStudy (anonymous):

I could of done that before -.- haha anyways so I get v1 - 2*v2=0 and 0=0 does that mean v2 is a free variable?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

Ok now Im confused on how to write the vector....

OpenStudy (amistre64):

2/4-3 = 2/1 L=3, Ev = <2,1>

OpenStudy (amistre64):

solve for v1 in terms of v2 and solve v2 in terms of v2

OpenStudy (anonymous):

oh nevermind I get it!!

OpenStudy (anonymous):

Thanks!!!

OpenStudy (amistre64):

yw